OpenStudy (anonymous):

a^3+b^3+c^3-3abc=0 Is their any identity related to it?

OpenStudy (lgbasallote):

no but there is \[a^3 + b^3 + c^3 + 3ab + 3ac + 3bc = 0\]

OpenStudy (lgbasallote):

for -3abc i believe there isnt any

OpenStudy (anonymous):

i know there is some identity including -3abc

OpenStudy (lgbasallote):

well that's the closest i can think of. goodluck finding -3abc hehe

OpenStudy (anonymous):

got it, your luck favoured me :D (a+b+c)(a^2+b^2+c^2-ab-bc-ca)

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

i think it is \[(a+b+c)(a^2+b^2+c^2-ab-bc-ca)=a^3+b^3+c^3-3abc\]

OpenStudy (lgbasallote):

hah told you it works "[

OpenStudy (lgbasallote):

:P *